# Modeling spectral filtering effects on color-matching functions: Implications for observer variability

### v1

@@ 4.4.1 Experiment and optimization method
The experimental scope of this study presents a notable limitation, as it involved only a single observer performing a color-matching experiment to derive color-matching functions (CMFs) with and without a physical filter. This constrained setup was sufficient to validate the method’s ability to estimate a spectral filter that transforms between the two sets of CMFs, successfully recovering the transmittance properties of the physical filter to some extent. However, it remains uncertain whether the approach would perform similarly for other observers. The lack of diversity in the observer pool limits the generalizability of the findings and raises questions about the robustness of the methodology across individuals. Future experiments involving a broader range of observers would be essential to evaluate whether the method consistently captures the physical properties of filters across different observers.
@@
As we saw in the experimental section, our optimization method works rather well for our purposes. However, there are other algorithms in this space which are relevant here. In [17] Finlayson and Zhu present an alternating least-squares (ALS) algorithm which also incorporates a smoothness constraint. The ALS algorithm works iteratively. First the best filter is found (for a given initial linear transform) then this is held fixed and the best linear transform is solved for. The linear transform is then held fixed and a new filter is found. The algorithm iterates like this until convergence. The ALS approach has the advantage that for a given initialisation condition it will always minimize the fitting error though the process may not necessarily converge to a global optimum. However, also in [17] it is shown how multiple initialisation conditions might be used to find the best overall result. Relatedly, in [20] the optimal transform and filter is found by warping the chromaticity spaces of the two sensor sets. Here a global optimum is found.
In the literature, the method of Riverz [21] is the most similar to Equations (2) through (10). There, his simplified method has a similar formulation to the one we derived above, it too is a null-space method. Rivertz applies his null-space method to the problem of making camera sensors colorimetric (a filter and linear transform is found that together make a particular camera better able to make colorimetric measurements). However, that study found that the ALS method worked better for a mean-square error metric. Rivertz’s null-space method was found be a more competitive optimisation when the Vora-Value[22] metric is used to assess similarity to the human visual system’s colour matching functions. With respect to the Vora-Value, the ALS and Rivertz’s simplified method delivered similar levels of performance. However, the ALS method was recently modified so it could optimise for the Vora Value [16]. Here, ALS always works better than the null-space method.
@@ 4.4.2 Modeling inter-observer variability
The current approach assumes that differences between CMF datasets can be attributed primarily to optical filtering effects, such as lens yellowing, while other physiological factors remain unmodeled. For instance, variations in macular pigment density, cone photopigment optical densities, and spectral shifts in cone photopigment sensitivity—all parameters described in Stockman’s cone fundamentals[13, 23]—are known to influence color perception but are not explicitly incorporated into this framework. These factors could interact with optical filtering effects in complex ways that are not captured by a single spectral filter and linear transformation matrix. Future work should aim to integrate these physiological parameters into the model to provide a more comprehensive understanding of inter-observer variability in CMFs and disentangle the contributions of different factors to observed dataset differences.
@@ 5 Conclusion
Future work should focus on incorporating a wider range of physiological parameters in the existing model for a more complete picture of inter-observer variability. Additionally, increasing the observer pool and refining experimental and optimization methods will enhance generalizability and address limitations observed in single-observer studies. These efforts will provide deeper insights into individual variability in color perception and further strengthen the framework developed in this study.
Lastly, building on the apparatus detailed in our previous publication[19], this study demonstrates that despite experimental noise and changes in conditions—such as the addition of a filter in front of the bipartite field—the observer was able to perform reliable color-matching. The results effectively captured the impact of the spectral filter on color-matching behavior, highlighting the stability of the apparatus in accommodating such modifications. While noise was present, the experimental setup proved robust enough to facilitate meaningful measurements, supporting its continued use in color vision studies.

### v2

@@ 4.4.1 Experiment and optimization method
The experimental scope of this study presents a notable limitation, as it involved only two observers performing a color-matching experiment to derive color-matching functions (CMFs) with and without a physical filter. Given the demographic differences between the two observers, this constrained setup was sufficient to validate the method’s ability to estimate a spectral filter that transforms between the two sets of CMFs, successfully recovering the transmittance properties of the physical filter to some extent. However, it remains uncertain whether the approach would perform similarly for other observers. The limited amount of observers limits the generalizability of the findings and raises questions about the robustness of the methodology across individuals. Future experiments involving a broader range of observers would be essential to evaluate whether the method consistently captures the physical properties of filters across different observers.
@@
As we saw in the experimental section, our optimization method works rather well for our purposes. However, there are other algorithms in this space which are relevant here. In [FinlaysonZhu] Finlayson and Zhu present an alternating least-squares (ALS) algorithm which also incorporates a smoothness constraint. The ALS algorithm works iteratively. First the best filter is found (for a given initial linear transform) then this is held fixed and the best linear transform is solved for. The linear transform is then held fixed and a new filter is found. The algorithm iterates like this until convergence. The ALS approach has the advantage that for a given initialisation condition it will always minimize the fitting error though the process may not necessarily converge to a global optimum. However, also in [FinlaysonZhu] it is shown how multiple initialisation conditions might be used to find the best overall result. Relatedly, in [FinlaysonFarup] the optimal transform and filter is found by warping the chromaticity spaces of the two sensor sets. Here a global optimum is found.
In the literature, the method of Riverz [Rivertz] is the most similar to Equations (2) through (10). There, his simplified method has a similar formulation to the one we derived above, it too is a null-space method. Rivertz applies his null-space method to the problem of making camera sensors colorimetric (a filter and linear transform is found that together make a particular camera better able to make colorimetric measurements). However, that study found that the ALS method worked better for a mean-square error metric. Rivertz’s null-space method was found be a more competitive optimisation when the Vora-Value[Vora] metric is used to assess similarity to the human visual system’s colour matching functions. With respect to the Vora-Value, the ALS and Rivertz’s simplified method delivered similar levels of performance. However, the ALS method was recently modified so it could optimise for the Vora Value [FinlaysonZhuVora]. Here, ALS always works better than the null-space method.
@@ 4.4.2 Modeling inter-observer variability
The current approach assumes that differences between CMF datasets can be attributed primarily to optical filtering effects, such as lens yellowing, while other physiological factors remain unmodeled. For instance, variations in macular pigment density, cone photopigment optical densities, and spectral shifts in cone photopigment sensitivity—all parameters described in Stockman’s cone fundamentals[cie2006fundamental, stockman2000spectral]—are known to influence color perception but are not explicitly incorporated into this framework. These factors could interact with optical filtering effects in complex ways that are not captured by a single spectral filter and linear transformation matrix. Future work should aim to integrate these physiological parameters into the model to provide a more comprehensive understanding of inter-observer variability in CMFs and disentangle the contributions of different factors to observed dataset differences.
@@ 5 Conclusion
Future work should focus on incorporating a wider range of physiological parameters in the existing model for a more complete picture of inter-observer variability. Additionally, increasing the observer pool and refining experimental and optimization methods will enhance generalizability to the broader population. These efforts will provide deeper insights into individual variability in color perception and further strengthen the framework developed in this study.
Lastly, building on the apparatus detailed in our previous publication[ragoo2024apparatus], this study demonstrates that despite experimental noise and changes in conditions—such as the addition of a filter in front of the bipartite field—the observer was able to perform reliable color-matching. The results effectively captured the impact of the spectral filter on color-matching behavior, highlighting the stability of the apparatus in accommodating such modifications. While noise was present, the experimental setup proved robust enough to facilitate meaningful measurements, supporting its continued use in color vision studies.